s Law for this (F--ks): Where F is the springs restoring force; k is the spring constant; and s is the stretch. The negative sign means the springs restoring force is opposite the stretch direction. You have a plot from weight IN] versus stretch [m). The data forms a linear trend y 3.662x+1,67. How much will the spring stretch if 51.7 grams is hung on the spring? Answer in centimeters with three significant figures or N/A if not enough information is given to answer. When you calculate your ansswer, dont use the negative sign in the Hookes Law formula. Just know that the negative sign simply denotes the force direction is opposite the stretch (or compression) Answer
Answer #1

using F = -k x = m g

given that Y = 3.662 x + 1.67

we know tha Y = mg  

so

3.662 x + 1.67 = m g

(3.662 x + 1.67) = 0.0517 * 9.81

x = -0.317 m